OpenStudy (anonymous):

Sean is deciding whether to select a satellite receiver or cable for his television programming. The satellite receiver costs $298.90 and the monthly charge is $68.70. With cable, there is no initial cost to purchase equipment, but the monthly charge for comparable channels is $74.80. After how many months will the total cost of the two systems be equal?

OpenStudy (anonymous):

49 months.

OpenStudy (anonymous):

How did you get that josh?

OpenStudy (anonymous):

298.9+68.7=74.8 6.1=289.9 49

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

Total cost equal: 74.8m = 298.9+68.7m -> 74.8m - 68.7m = $289.9 Number of months: => m = $ 289.9 / 6.1 = ~29 months

OpenStudy (anonymous):

I dont think Chlorophyll is correct. Use A for the cost after, and B for months. A=289.9+68.7B A=68.7 289.9+38.7b=74.8b 6.1b=298.9 b=49